i am having trouble understanding these types of questions..can you help?
length of the major axis of the ellipse 3x^2+2y^2=6 is __________?
this is a problem we had in our class to be reviewed tomorrow..thanks.

Answer by stanbon(75887)

length of the major axis of the ellipse 3x^2+2y^2=6 is __________?
------------
Divide thru by 6 to see it in standard form:
x^2/2 + y^2/3 = 1
The larger denominator determines the "major" axis.
For you it is a^2=3, so a=sqrt3, and 2a=2sqrt3
So the lenth of the major axis is 2sqrt3
--------------
Cheers,
Stan H.